A restaurent has a total of 27 tables and booths.Each table seats up to 8 people and each booth seats upto 4 people.The restaurant can seat upto a total of 140 people.How many booths are in restaurant?
Answer by fractalier(6550)

Let the numbers of tables and booths be t and b. Then we write
t + b = 27 and
8t + 4b = 140
Now multiply the top equation by 4 and subtract
8t + 4b = 140
-(4t + 4b = 108)
-----------------
4t = 32
t = 8 tables
which means there are
b = 19 booths